OpenStudy (anonymous):

Simplify (200-100)/4+5

OpenStudy (anonymous):

100/9 is 11.1111111111

OpenStudy (anonymous):

(200-100)/4+5 or (200-100) --------- 4+5 in the first one the answer is 100/4+5=25+5=30 in the second one its 100/9

OpenStudy (anonymous):

nope the answer is 100/4=25+5=30